# Disentangling $\gamma-\beta$: the 4th-order velocity moments based on spherical Jeans analysis

### v1

@@ 3.1 General comparison between 2nd-order only and 2nd & 4th-order fitting
3.2 Effects of varying the number of samples stars
The impact of varying the number of sample stars is shown in Figure 4. We perform fittings by randomly selecting 500 stars (the left panels) and 5,000 stars (the right panels) from Mock Galaxy 01 (the upper row) and assigning $\Delta v_{\rm los}=2$ km s-1. Figure 4 reveals that changing the number of sample stars mainly affects the tightness of uncertainty of the free parameters’ posterior distribution. This feature is more pronounced in $\beta$ where its uncertainty is shown to be reduced by a factor of $\sim$3. Notably, a similar trend of a substantial reduction in $\beta$ uncertainty is observed when we fit 500 and 5,000 sample stars from Mock Galaxy 02 (the lower row), also with $\Delta v_{\rm los}=2.0$ km s-1. However, the lower row of Figure 4 shows a strong systematic bias for both $\beta$ and $\gamma$. The inferred $\beta$ is biased toward the more tangential regime, and $\gamma$ is biased toward a cuspier profile than the input value. Increasing the sample size to $N=$ 5,000 stars is helpless because it only reduces the random error while maintaining very similar median values, indicating that the systematic bias remains unimproved. These results suggest that it is not the number of sample stars holding the most control over this systematic bias. We explain the reason behind this bias in the next section.
Figure 4: The same as Figure 3 but in the left panels we use 500 sample stars, while in the right panels we use 5,000 sample stars. The upper (lower) row shows the results for Mock Galaxy 01 (02). All panels are obtained from fitting using the 2nd & 4th-order moments.
3.3 Effects of varying the l.o.s. velocity measurement error
Figure 5 shows the marginalization of free parameters with different $\Delta v_{\rm los}$ for Mock Galaxy 02. In the left panel, we assign $\Delta v_{\rm los}=2.0$ km s-1 while in the right panel $\Delta v_{\rm los}=0.01$ km s-1. It is, of course, practically unachievable by the current instruments to have $\Delta v_{\rm los}=0.01$ km s-1. This tiny velocity error is selected in order to eliminate most of its effects to understand the model’s behavior related to velocity errors. Figure 6 is the same as Figure 5 but for Mock Galaxy 01.
We find a severe systematic bias for $\beta$ and $\gamma$ in the case of Mock Galaxy 02 with $\Delta v_{\rm los}=2.0$ km s-1 (the left panel of Figure 5). The systematic bias does not occur, however, when we push the velocity error to 0.01 km s-1 (the right panel of Figure 5). This velocity error-driven bias indicates the strong dependency of the 4th-order moments on the velocity distribution wings that have a small number of stars and, hence, are more severely affected by the velocity error. The departure from Gaussianity of the LOSVD, the feature that is tracked in 4th-order moments fitting, is smeared out by the velocity error that has a Gaussian distribution.
Interestingly, the dramatic change of systematic bias does not occur in Mock Galaxy 01 fitting. Figure 6 shows that varying $\Delta v_{\rm los}$ from 2.0 km s-1 to 0.01 km s-1 does not provide a considerably better result. These results from Mock Galaxy 01 and Mock Galaxy 02 fitting suggest that it is not the absolute value of $\Delta v_{\rm los}$ that matters, but the ratio of $\sigma_{\rm los,global}/\Delta v_{\rm los}$ instead. As long as $\sigma_{\rm los,global}\approx\Delta v_{\rm los}$, like the case of Mock Galaxy 02 fitting with $\Delta v_{\rm los}=2$ km s-1, the result will most likely be biased. However, after the ratio reaches $\sigma_{\rm los,global}/\Delta v_{\rm los}\approx 4$ (like the case of Mock Galaxy 01 fitting), not only the systematic bias due to the velocity error is not observed anymore, but also there is no significant improvement if we push the ratio to be larger.
To have a further test regarding this characteristic and check if the velocity error-driven bias can be addressed by increasing the number of stars, we perform a similar fitting on Mock Galaxy 01 and randomly select $N=10^{4}$ stars. We assign $\sigma_{\rm los,global}/\Delta v_{\rm los}=1.7$, which is very similar to the case of Mock Galaxy 02 fitting with $\Delta v_{\rm los}=2$ km s-1. We find that a very similar bias, where the inferred $\beta$ is far more tangential and $\gamma$ is cuspier than the true value, is observed. We conclude that this velocity error-driven bias cannot be reduced by increasing the sample size. Fortunately, among all the fitting, $\beta$ is consistently recovered as long as $\sigma_{\rm los,global}/\Delta v_{\rm los}\gtrsim 4$ even though the number of sample stars is only 500 and $\gamma$ is inconclusive.
[…]
Figure 6: The same as Figure 5 but for Mock Galaxy 01.
3.4 Effects of varying the $R_{\rm max}$ of stars
We investigate the role of stars’ spatial distribution by conducting some fittings using different values for the furthest member star available $R_{\rm max}$, while fixing the Plummer scale-length $a_{*}$ as well as the dark matter scale-length $a_{\rm dm}$, with $b=a_{\rm dm}/a_{*}=10$. We set four different hard cuts $R_{\rm max}=[\textrm{no limit,}~{}a_{\rm dm},~{}0.5a_{\rm dm},~{}0.1a_{\rm dm}]$ and randomly select 5,000 member stars from Mock Galaxy 02 for each $R_{\rm max}$. In order not to mix with the effect from the velocity error, we fix $\Delta v_{\rm los}=0.01$ km s-1.
Results of the fitting are shown in Figure 7 where it is $R_{\rm max}=\textrm{no limit}$, $R_{\rm max}=a_{\rm dm}$, $R_{\rm max}=0.5a_{\rm dm}$, and $R_{\rm max}=0.1a_{\rm dm}$, respectively from left to right. Figure 7 illustrates that stars in the inner region play a dominant role in constraining $\gamma$, as evidenced by the nearly indistinguishable results among the first three leftmost panels. Moreover, the median values of $\gamma$ for limited $R_{\rm max}$ show slightly improved performance compared to the case with no limit on $R_{\rm max}$. This slight improvement is likely due to the higher number of stars in the inner part of the galaxy probing $\gamma$, a behavior that is also observed in Chang & Necib (2021). Only when $R_{\rm max}=0.1a_{\rm dm}=a_{*}$ the posterior of $\gamma$ shows some noticeable difference with a larger offset of the median value from the true value and a larger random error compared to the other three cases. This indicates that even though stars in the inner part of the galaxy play the main role in estimating $\gamma$, an extensive sample of at least up to $\sim 0.5a_{\rm dm}$ is still necessary. However, $a_{\rm dm}$ is not observable, and we rely only on $a_{*}$ to evaluate the extent of a sample set. Because of its dissipation nature, baryonic matter tends to be more centrally concentrated than dark matter (White & Rees, 1978; Efstathiou & Silk, 1983). It is thus important to assemble sample stars beyond $R=a_{*}$ to place useful limits on $\gamma$
[…]
Figure 7 also shows that the estimation of $\beta$ is consistent across all scenarios. There is no noticeable effect from changing $R_{\rm max}$ nor the decreasing quality on the recovery of $\gamma$ when $R_{\rm max}=0.1a_{\rm dm}$. This $\beta$ and $\gamma$ independency confirm the reduced degeneracy between the two parameters and different reliance on the observational properties after 4th-order moments are incorporated.
4 Discussion
4.1 Dependency on the prior range
Jeans modeling has a limitation in that the solutions it provides are not guaranteed to be physically plausible. However, the prior range of $\gamma$ has to be carefully selected to ensure that the results do not carry any effect from the choice. Therefore, while negative $\gamma$ is physically unlikely, it is incautious if some range in $\gamma<0$ is not included in the prior because it might suffer from the border effect. An unduly restrictive prior does not give enough solution space around $\gamma=0$ and will force the walkers to explore $\gamma>0$. The cored profile is thus systematically disfavoured.
We also examine if this issue has any effects when the input density profile is a cuspy halo. The result is shown in Figure 8 where we fit 5,000 randomly selected stars from Mock Galaxy 01 with two different ranges of prior on $\gamma$, i.e., a broad range $-2.0\leqslant\gamma\leqslant 2.0$, and a narrower range of $0\leqslant\gamma\leqslant 2.0$ that exclude the unphysical region. The shift of $\gamma$’s posterior shown in Figure 8 suggests that the issue also affects the distribution even for a cuspy true density profile. For a cuspy profile, this should not occur if the model, with the provided data, can completely exclude $\gamma<0$. To avoid this prior-caused bias on $\gamma$, we decide to apply the broad range of prior for $\gamma$, $-2.0\leqslant\gamma\leqslant 2.0$, for all the fittings. Prior dependency in the estimation on $\gamma$ is also observed in Read et al. (2018), Genina et al. (2020), Chang & Necib (2021), and Hayashi et al. (2023b), especially for a smaller sample size.
Figure 8: The same as Figure 4 but the left panel is fitting using $-2.0\leqslant\gamma\leqslant 2.0$ and the right panel is fitting using $0\leqslant\gamma\leqslant 2.0$ for Mock Galaxy 01. Both panels are obtained from fitting using 2nd & 4th-order moments.
4.2 The recovery of $\gamma$
As one of the central focuses of this work, in this subsection we discuss the recovery of $\gamma$. It is interesting to note that the case for a cored halo density profile can be recovered with a robust exclusion of the cuspy one. However, it does not occur in the cuspy case where the posterior distribution of $\gamma$ does not peak at $\gamma\approx 1$ as well as there is considerable probability in the cored region. This difference can be explained by the fact that $\kappa_{\rm los}(R)$ deviates from Gaussianity more strongly, especially in the inner part. This strong deviation gives more signatures for the 4th-order moments to exclude the incorrect regions.
Figure 9 shows the change of $\kappa_{\rm los}(R)$ for different values of $b\equiv a_{\rm dm}/a_{\ast}$, i.e., the ratio between the scale lengths of dark and luminous matter, in a specific $\gamma$. From left to right, the values of $\gamma$ are $[0.0,1.0,1.5,2.0]$, respectively. In each panel, each line denotes $b=[1,5,10]$, respectively, from the lightest to the darkest, while the other parameters are fixed. It is shown that in the case of the cuspier density profile, the range of $\kappa_{\rm los}(R)$ is narrower, being more concentrated around $\kappa_{\rm los}\approx 3$. On the other hand, the range of $\kappa_{\rm los}(R)$ for a cored density profile is broader, reaching regions that are unlikely to be reached by a cuspier profile. Therefore, a cored halo has more opportunities to be identified more robustly than a cuspy one. The narrow range of $\kappa_{\rm los}(R)$ in a cuspy halo also explains the more rapid decrease in all the posterior of $\gamma$ for region $\gamma>1$ compared to $\gamma<0$. The crowded region around $\kappa_{\rm los}\approx 3$ explains the reason for the difficulty in Mock Galaxy 01 fitting to rule out $\gamma\approx 0$. Another feature to note in Figure 9 is the difference of $\kappa_{\rm los}(R)$ between a core and a cusp with different values of $b$. The gap in $\kappa_{\rm los}(R)$ between the cored and cuspy profile is more pronounced for a larger $b$, allowing a halo with a large $b$ to be distinguished more easily.
Why does a cored profile obtain a more significant $\kappa_{\rm los}(R)$ boost when $b$ increases, compared to a cuspy profile? This can be explained by, firstly, understanding that the non-Gaussian LOSVD in a system with $\beta=0$ is a consequence of projection, i.e., integrating velocity distribution with different values of dispersion and star count. The smallest galactocentric distance along a certain $R$, hence $r_{\rm min}=R$, has the most star count in that line-of-sight. Therefore, the projected velocity dispersion at that point $\sigma_{\parallel}(r_{\rm min})$ has the most influence in determining $\sigma_{\rm los}(R)$. Summing up velocity distributions along the l.o.s. with a lower star count but a larger dispersion to $\sigma_{\parallel}(r_{\rm min})$ will populate the velocity tails, resulting in a more heavy-tailed distribution than a Gaussian. The larger the ratio of $\sigma_{\parallel}(r_{\rm outer})/\sigma_{\parallel}(r_{\rm min})$, where $\sigma_{\parallel}(r_{\rm outer})$ is the outermost point a member star is bound to the galaxy, the heavier the velocity tails are and hence the larger $\kappa_{\rm los}$ at that line-of-sight. This means that if $\sigma_{\parallel}(r_{\rm outer})<\sigma_{\parallel}(r_{\rm min})$, then there is no region in that line-of-sight to populate the tails. Moreover, the addition of stars from $r>r_{\rm min}$ makes the tails less significant compared to the ’central’ part of the integrated velocity distribution. In this case, the resulting LOSVD has fewer outliers than a Gaussian.
Using $\sigma_{\parallel}(r_{\rm outer})/\sigma_{\parallel}(r_{\rm min})$ to evaluate if the resulting LOSVD has fewer or more outliers than Gaussian is, of course, a simplification and aims only to provide a general sense because $\sigma_{\parallel}(r)$ along the line-of-sight might wind up and down. The LOSVD is the net product of all contributions from each $r\geqslant r_{\rm min}$. To provide more illustration, let us consider galaxies having isotropic velocity anisotropy with, again, various $\gamma$ and $b$. Because of its isotropy, $\sigma_{r}(r)$ can represent any $\sigma_{\parallel}(r)$. Figure 10 shows normalized $\sigma_{r}$ as a function of galactocentric distance scaled to $a_{*}$. The values of $\gamma$ are $\gamma=[0.0,1.0,2.0]$, respectively, from left to right and $b=[10,5,1]$, respectively, from the most solid one in each panel.
Figure 10: Normalized velocity dispersion profile in $r$ direction for dark halo with various $\gamma$ and $b$. The values are $\gamma=[0.0,1.0,2.0]$, respectively, from left to right. Different shades of line solidness in each panel denote the difference in $b$ where $b=[10,5,1]$, respectively, from the most solid line.
The leftmost panel of Figure 10 shows that because of its relatively low central density, a cored profile always has a low velocity dispersion in the central part and generally increases with $r$, except when $b\lesssim 1$. This characteristic, on the other hand, does not exist in a cuspy profile where the velocity dispersion in the central part is relatively larger than a core and generally decreases with $r$. The low central velocity dispersion with a large number of stars in a core halo provides a window for stars in the outer region with higher velocity dispersion to be outliers when they are integrated along the line-of-sight to obtain the LOSVD. Furthermore, because the enclosed mass at a certain $r$ is larger in a dark halo with a larger $b$, the velocity dispersion also increases more rapidly. As a result, outliers are more significant in a halo with a larger $b$, in agreement with the simple concept of $\sigma_{\parallel}(r_{\rm outer})/\sigma_{\parallel}(r_{\rm min})$ ratio. This explains two characteristics; the first one is the reason for a considerable departure from $\kappa_{\rm los}(R)\approx 3$ for a cored dark halo with $b\gtrsim 5$ (the leftmost panel of Figure 9). The second is the reason for larger $\kappa_{\rm los}(R)$ everywhere for a cored dark halo compared to the cuspy one with the same $b$ (all panels in Figure 9).
Given the dependency of the estimation of $\gamma$ on $\kappa_{\rm los}(R)$, it is interesting to know if we can have an alert when the estimation is not representative or might suffer from systematic bias. However, unlike the mock data case, the intrinsic $\kappa_{\rm los}(R)$ is not accessible in the real condition. Fortunately, star distribution in dSphs and UFDs decreases rapidly beyond $a_{*}$. Hence, calculating the $\kappa_{\rm los}$ directly from the discrete kinematics data of observed stars is heavily influenced by $\kappa_{\rm los}$ in the inner region $R<a_{*}$. This is enough to provide a justification if the estimation of $\gamma$ might or might not be representative. In our sample sets, for example, the $\kappa_{\rm los}$ are 3.10, 3.31, 7.43, and 7.02, respectively, for Mock Galaxy 01 500 stars, Mock Galaxy 01 5,000 stars, Mock Galaxy 02 500 stars, and Mock Galaxy 02 5,000 stars. Even 500 sample stars are enough to provide a hint if $\kappa_{\rm los}(R)$ deviates considerably from Gaussianity in some of its parts.
The estimation of $\gamma$ is also challenged by a remaining degeneracy between parameters of dark matter distribution, i.e., $\gamma$, $b$, and $\rho_{0}$, which might be the nature of any parametric model. Figure 11 shows $\sigma_{\rm los}(R)$ response by varying either $b$ (the left panel), $\gamma$ (the middle panel), or $\rho_{0}$ (the right panel), while fixing the other two parameters. The values for each parameter are $b=[20,15,10,5]$, $\gamma=[1.5,1.0,0.5,0.0]$, and $\rho_{0}=[0.2,0.1,0.064,0.02]$ M pc-3, respectively, from the most solid line. When fixed, we select $b=10$, $\gamma=1.0$, and $\rho_{0}=0.064$ M pc-3. In Figure 12, the left panel shows the effect on $\kappa_{\rm los}(R)$ by varying $b=[10,5,3,1]$, respectively, from the most solid line while holding $\gamma=0.0$. In the right panel we change $\gamma=[0.0,0.5,1.0,1.5]$, respectively, from the most solid line while holding $b=10$. Note that $\kappa_{\rm los}(R)$ does not depend on $\rho_{0}$, leaving the estimation of $\rho_{0}$ completely relies on the 2nd-order moment $\sigma_{\rm los}(R)$ fitting. The similar response of $\sigma_{\rm los}(R)$ and $\kappa_{\rm los}(R)$ to changes in $b$, $\gamma$, or $\rho_{0}$ contributes to the difficulties in ruling out the incorrect values of $\gamma$, especially in the case of $\gamma=1.0$.
[…]
Figure 12: The line-of-sight kurtosis profile of dark halos with a fixed $\gamma=0.0$ but varying $b=[10,5,3,1]$ in the left panel, and a fixed $b=10$ but varying $\gamma=[0.0,0.5,1.0,1.5]$ in the right panel, respectively, from the most solid line.
@@ 4.3 The recovery of $\beta$
Figure 13 shows the behavior of $\kappa_{\rm los}(R)$ when $\beta$ changes. From the leftmost panel, the values of $\beta$ are -0.5, 0.0, and 0.5, respectively. Meanwhile, lines colored in orange, green, and pink denote different values of the dark matter inner density slope $\gamma=[0.0,1.0,2.0]$, respectively. We fix the ratio of dark matter scale length to light scale length as $b=10$ in all panels.
@@
From the previous section, we understand that we understand that increasing $\gamma$ results in more flat-topped LOSVDs with fewer outliers. This happens because the larger velocity dispersion pushes stars away from $v_{\rm los}\approx v_{\rm sys}$ in high stellar density regions. By contrast, as illustrated in Figure 13, increasing $\beta$ decreases $\kappa_{\rm los}(R)$ in the inner region but increases it in the outer region, according to the light and mass profiles adopted in our model. This different response of $\kappa_{\rm los}(R)$ to changes in $\beta$ and $\gamma$ is the main reason for the disentanglement in the estimation of the two parameters. Models relying solely on 2nd-order moments have no access to this distinction because increasing $\beta$ or $\gamma$ leads to an increase in $\sigma_{\rm los}(R)$, especially in $R<a_{\rm dm}$. Moreover, unlike a large $\gamma$ that produces a narrow range of $\kappa_{\rm los}(R)$ as shown in Figure 9, $\beta$ has no regime that confines $\kappa_{\rm los}(R)$ in a similar restriction. Even in the most subtle response of $\kappa_{\rm los}(R)$, that occurs when $\beta=0$, $\kappa_{\rm los}(R)$ is allowed to lie in a relatively wide range $2\lesssim\kappa_{\rm los}(R)\lesssim 6$ as shown in the middle panel in Figure 13. Thus, the model finds it less challenging to rule out incorrect values of $\beta$ than those of $\gamma$. Overall, $\kappa_{\rm los}(R)$ is more sensitive to changes in $\beta$ than to changes in $\gamma$, particularly in the case of large $\gamma$. This sensitivity is the main reason why applying a higher-order moment analysis is able to place tight constraints on $\beta$.
@@
Figure 13: The line-of-sight kurtosis profile of dark halos with a fixed $b=10$ but different values of $\beta$ and $\gamma$. The values are $\beta=[-0.5,0.0,0.5]$, respectively, from left to right. Orange, green, and pink lines denote $\gamma=[0.0,1.0,2.0]$, respectively.
@@
The common trend of increasing $\kappa_{\rm los}(R)$ with increasing $\beta$ in the outer region $R\gtrsim a_{\rm dm}\approx 10a_{*}$ can be explained by considering a line-of-sight in that region. The stellar density in this region drops rapidly as $\nu(r)\propto r^{-5}$. Therefore, any l.o.s. will be highly dominated by the stars located around the closest point of the l.o.s. to the galactic center $r_{\rm min}=R$. In a system with $\beta>0$, this point is filled by stars with $v_{\rm los}\approx v_{\rm sys}$, resulting in a strong-peaked LOSVD around $v_{\rm los}=v_{\rm sys}$ and, hence, a large $\kappa_{\rm los}$. Meanwhile, for a system with $\beta<0$, this point is filled by a diverse value of $v_{\parallel}$ from stars with nearly circular orbits with different inclinations, resulting in a broad and flat-topped LOSVD. A similar behavior is mentioned by Binney & Tremaine (2008) for a Hernquist density model for both light and mass profiles. In addition, along any line-of-sight, the velocity ellipsoid component parallel to the line of sight, $\sigma_{\parallel}(r)$, is always more aligned to $\sigma_{r}(r)$ with increasing $r$. Because of this reason, high $v_{\rm los}$ stars in $r\gg r_{\rm min}$ populate the LOSVD tails in a system with radially biased orbits. However, it does not occur in a system with tangentially biased orbits because the l.o.s. cuts stars’ orbits at approximately right angles.
@@
The behavior of $\kappa_{\rm los}(R)$ in the inner region is more complex due to the interplay between the dark matter density profile and stellar density. Instead of being heavily dominated by stars in $r_{\rm min}$, the integrated LOSVD is more dominated by stars lying in front of or behind $r_{\rm min}$ because of the shallow stellar density gradient. Because of that, contribution to $v_{\rm los}$ in these regions is mainly obtained from the radial component of the star’s orbital velocity. If $\beta<0$, most of the stars in those regions have $v_{\rm los}\approx v_{\rm sys}$, resulting in a strong spike around $v_{\rm los}=v_{\rm sys}$ in the LOSVD and a large $\kappa_{\rm los}$. If $\beta>0$, the LOSVD is filled by stars with various $v_{\rm los}$ from the radial velocity of stars with different galactocentric distances. This effect produces a broad, flat-topped LOSVD and a small $\kappa_{\rm los}$.
The more sensitive response of $\kappa_{\rm los}(R)$ in a system with $\beta\neq 0$ suggests that the estimation of free parameters can be achieved more robustly than in $\beta=0$ case. However, using different mass modeling methods, Read et al. (2021) find that it is more challenging to recover the density and velocity anisotropy in systems with $\beta>0$ than systems with $\beta\leqslant 0$. Merrifield & Kent (1990) also suggest that the number of stars required to recover kurtosis in a heavy-tailed distribution is significantly larger than in a more thin-tailed distribution. One more thing to note is our assumption that $\beta(r)=$ constant. In reality, $\beta$ might change with $r$ as suggested by various studies. For example, Loebman et al. (2018) simulate Milky Way-sized galaxies and observed a tendency for $\beta\sim 0$ in the innermost regions, changing to a radially biased $\beta\sim 0.7$ in the outer regions. Meanwhile, Kowalczyk et al. (2019) apply a spherically symmetric Schwarzschild orbit superposition method to the Fornax dSph. Using this method, which does not require any assumption on the shape of $\beta(r)$, they find $\beta\approx 0$ in the center and mildly decreases with radius. Therefore, a careful study of systems with $\beta\neq 0$ as well as non-constant $\beta$ is necessary.

### v2

@@ 4.1.2 Degeneracy between $\gamma-\rho_{0}-b$
4.1.3 Effects of varying the $R_{\rm max}$ of stars
We examine the impact of the spatial distribution of stars by performing fittings with varying values of the maximum radial distance of the outermost member star, $R_{\rm max}$, while keeping the Plummer scale-length, $a_{*}$, and the dark matter scale-length, $a_{\rm dm}$, fixed. The ratio of these two scales is set to $b=a_{\rm dm}/a_{*}=10$ for the cuspy dSph and $b=4.5$ for the cored dSph mock galaxies. Two hard cuts are applied: $R_{\rm max}=100{\rm~{}pc}=a_{*}$ and $R_{\rm max}=500$ pc. The latter is selected to approximate the typical tidal radius of Milky Way dSph satellites, which is typically a few times the half-light radius. For each dSph mock galaxy, we generate 10 realizations by randomly selecting 5,000 member stars and assigning a negligible l.o.s. velocity error of $\delta v_{\rm los}=0.01$ km s-1 to isolate the spatial limitation effects.
Results of the fitting are shown in Figure 10 where green indicates $R_{\rm max}=100$ pc and violet indicates $R_{\rm max}=500$ pc. In the left panel, we show the density profile recoveries in the range of $1$ $10^{4}$ pc. The range in this panel lies well beyond the availability of tracers in order to show the hard-cut effects of $R_{\rm max}$. The black solid line denotes the true profile, and the darker (lighter) shade denotes the 68% (95%) confidence level. In the right panel, we show the posterior distributions for $\gamma$ and $\beta$ with the same color code. Red lines denote the true values and three vertical dashed lines denote the median and 95% confidence level. Meanwhile, contour lines in the 2-dimensional posterior distributions represent the 68% and 95% confidence levels.
The results show that the true density profile is recovered within the 68% confidence level in all scenarios. However, there are increasing uncertainties and shifts in the median value as $R_{\rm max}$ is more restrictive. This occurs not only in the outer region, as expected, but also in the inner region. Moreover, contours for $R_{\rm max}=100$ pc reveal signs of a stronger degeneracy between $\gamma$ and $\beta$, suggesting that the spatial extent of tracers contributes to breaking the mass-anisotropy degeneracy. This indicates that even though stars in the inner region play a dominant role in capturing the overall density profile trends, an extensive sample of up to a few $a_{*}$ is important to place useful limits on $\gamma$.
Figure 10: Density profile recovery and posterior distributions for $\gamma$ and $\beta$ obtained from the fittings of 5,000 stars from the cuspy dSph mock galaxy and setting hard cuts on the tracer’s spatial distribution at $R=5a_{*}=500$ pc (violet) and $R=a_{*}=100$ pc (green). All results are obtained by applying the model that incorporates 4th-order moments. In the left panel, the black solid line indicates the true profile and the darker (lighter) shade denotes the 68% (95%) confidence level. In the right panel, red solid lines denote the true values of each parameter, while the three vertical dashed lines indicate the median and 95% confidence intervals. Contour lines in the 2-dimensional posterior distributions represent the 68% and 95% confidence levels.
@@ 4.2 The recovery of $\beta$
Apart from the estimation of $\gamma$, the estimation of $\beta$ is shown to be undisturbed by that situation of whether or not incorrect $\gamma$ can be ruled out. As long as $\sigma_{\rm los,global}/\delta v_{\rm los}\gtrsim 4$ is satisfied and the sample correctly traces the intrinsic kurtosis profile, $\beta$ is shown to be constrained within 95% confidence level. To investigate the main reason for the disentanglement between $\gamma$ and $\beta$, we provide Figure 11 that shows the behavior of l.o.s. kurtosis profile when $\beta$ changes. The leftmost panel corresponds to a system with tangentially biased orbits having $\beta=-0.5$, $\beta=0.0$ for the middle panel, and $\beta=0.5$ for the right panel. From the leftmost panel, the values of $\beta$ are $-0.5$, 0.0, and 0.5, respectively. Lines colored in orange, green, and pink denote different values of the dark matter inner density slope $\gamma=[0.0,1.0,2.0]$, respectively. We fix the ratio of dark matter scale length to light scale length as $b=10$ in all panels.
@@
From the previous section, we understand that increasing $\gamma$ results in smaller $\kappa_{\rm los}(R)$, which means more flat-topped LOSVDs with fewer outliers. By contrast, as illustrated in Figure 11, an increase in $\beta$ leads to a decrease in $\kappa_{\rm los}(R)$ in the inner region while causing an increase in the outer region, according to the light and mass profiles adopted in our model. This different response of $\kappa_{\rm los}(R)$ to changes in $\beta$ and $\gamma$ is the main reason for the disentanglement in the estimation of the two parameters. Models relying solely on 2nd-order moments have no access to this distinction because increasing $\beta$ or $\gamma$ leads to a hardly distinguished increase in $\sigma_{\rm los}(R)$, particularly in $R<a_{\rm dm}$. Moreover, unlike a large $\gamma$ that produces a narrow range of $\kappa_{\rm los}(R)$ as shown in Figure 7, $\beta$ has no regime that confines $\kappa_{\rm los}(R)$ in a similar restriction. Even in the most subtle response of $\kappa_{\rm los}(R)$, that occurs when $\beta=0$, $\kappa_{\rm los}(R)$ is allowed to lie in a relatively wide range $2\lesssim\kappa_{\rm los}(R)\lesssim 6$ as shown in the middle panel in Figure 11. Thus, the model finds it less challenging to rule out incorrect values of $\beta$ than those of $\gamma$. This sensitivity is the main reason why applying a higher-order moment analysis is able to place tight constraints on $\beta$.
@@
Figure 11: The line-of-sight kurtosis profile of dark halos with a fixed $b=10$ but different values of $\beta$ and $\gamma$. The values are $\beta=[-0.5,0.0,0.5]$, respectively, from left to right. Solid orange, dashed green, and dash-dotted pink lines denote $\gamma=[0.0,1.0,2.0]$, respectively.
@@
In Figure 11, $\kappa_{\rm los}$ profiles in the outer region ($R\gg a_{*}$) monotonically increase as $\beta$ increases. This common trend can be explained by considering a line-of-sight in ($R\gg a_{*}$) region. The stellar density in this l.o.s. drops rapidly as $\nu(r)\propto r^{-5}$. Therefore, the l.o.s. will be highly dominated by the stars located around the closest point of the l.o.s. to the galactic center $r_{\rm min}=R$. In a system with $\beta>0$, this point is filled by stars with $v_{\rm los}\approx v_{\rm sys}$ because the l.o.s. cuts stellar orbits almost perpendicularly, resulting in a strong-peaked LOSVD around $v_{\rm los}=v_{\rm sys}$ and, hence, a large $\kappa_{\rm los}$. Meanwhile, for a system with $\beta<0$, this point is filled by a diverse value of $v_{\parallel}$ from stars of nearly circular orbits with different inclinations, resulting in a broad and flat-topped LOSVD. A similar behavior is mentioned by Binney & Tremaine (2008) for a Hernquist density model for both light and mass profiles. In addition, along any l.o.s., the velocity ellipsoid component parallel to the line of sight, $\sigma_{\parallel}(r)$, is always more aligned to $\sigma_{r}(r)$ with increasing $r$. Because of this reason, high $v_{\rm los}$ stars in $r\gg r_{\rm min}$ populate the LOSVD tails in a system with radially biased orbits. However, it does not occur in a system with tangentially biased orbits because the l.o.s. cuts stars’ orbits at approximately right angles.
@@
The behavior of $\kappa_{\rm los}(R)$ in the inner region is more complex due to the more various interplay between the dark matter density profile and stellar density profile. Instead of being heavily dominated by stars at $r_{\rm min}$, the integrated LOSVD is more dominated by stars lying in front of or behind $r_{\rm min}$ within the core Plummer profile due to the shallow stellar density gradient in that region. Because of that, contribution to $v_{\rm los}$ in these regions is mainly obtained from the radial component of the star’s orbital velocity. If $\beta<0$, most of the stars in those regions have $v_{\rm los}\approx v_{\rm sys}$, resulting in a strong spike around $v_{\rm los}=v_{\rm sys}$ in the LOSVD and a large $\kappa_{\rm los}$. If $\beta>0$, the LOSVD is filled by stars with various $v_{\rm los}$ from the radial velocity of stars with different galactocentric distances. This effect produces a broad, flat-topped LOSVD and a small $\kappa_{\rm los}$.
Even though the observed sensitivity of $\kappa_{\rm los}(R)$ in systems where $\beta\neq 0$ indicates a potential tendency for more robust parameter estimation compared to the $\beta=0$ case, further investigation is still needed to confirm whether this sensitivity consistently leads to improved robustness in parameter estimation. Merrifield & Kent (1990) suggests that the number of stars required to recover kurtosis in a heavy-tailed distribution is significantly larger than in a more thin-tailed distribution. This can intuitively be understood because to confirm that the distribution tails are strong, there must be a sufficient number of stars populating the relatively low probability regions. To conclude a thin-tailed distribution, by contrast, the stars just need to be absent. This is confirmed by Read et al. (2021) that apply four mass modeling methods to mock data and find that it is more challenging to recover the density profile and velocity anisotropy parameter in a system with $\beta(r)$ follows an Osipkov-Merrit profile (Osipkov, 1979; Merritt, 1985) whose $\kappa_{\rm los}(R)>3$ across the entire range of radii.
It is also important to consider our assumption that $\beta(r)$ is constant. In reality, $\beta$ may vary with radius, as suggested by various studies. For example, Loebman et al. (2018) simulate Milky Way-sized galaxies and observed a tendency that $\beta\sim 0$ in the innermost regions, transitioning to a radially biased $\beta\sim 0.7$ in the outer regions. Meanwhile, Kowalczyk et al. (2019) analyzes the Fornax dSph using a spherically symmetric Schwarzschild orbit superposition method, which does not require any assumption on the shape of $\beta(r)$. They find $\beta\approx 0$ near the center but gradually decrease with radius. These varied findings highlight the necessity of carefully studying systems with $\beta\neq 0$, as well as those with radially varying $\beta$ to obtain more accurate insight.